Which of the following mixtures may be act as a buffer solution?

A) HBr, NaBr

B) HF, NaF

C) HCl, NaCl

D) HI, NaI
Answer #1
Concepts and reason

The concept used in this problem is the buffer system and the concept of conjugate acid-base pair.

Initially, the mixtures which do not act buffer solution is identified. After that, the mixture which acts as a buffer solution is identified.

Fundamentals

Buffer solution:

Buffer is the solution which keeps it's pH constant even on the addition of a small amount of acids or base. The buffer is used to maintain the pH of the reaction. The buffers are two types: acidic buffer and basic buffer.

Acidic buffer:

The buffer system contains a weak acid and its conjugate base is called an acidic buffer. The pH of the acidic buffer is below 77 . The example of an acidic buffer is a mixture of acetic acid and sodium acetate.

Basic buffer:

The buffer system contains a mixture of a weak base and its conjugate acid is called basic or alkaline buffer. The pH of the basic buffer is above 77 . The solution of ammonium hydroxide and ammonium chloride is called a basic buffer.

The hydrogen bromide (HBr)\left( {{\rm{HBr}}} \right) is a strong acid because the size of bromide ion is very big and the length of the bond between hydrogen and bromide ion is very large which causes weak bond strength. It dissociates completely to form hydrogen ion and bromide ion and equilibrium cannot be established. Hence, the mixture of HBr{\rm{HBr}} and NaBr{\rm{NaBr}} cannot acts as a buffer.

HBr+NaOHNaBr+H2O{\rm{HBr}} + {\rm{NaOH}} \to {\rm{NaBr}} + {{\rm{H}}_2}{\rm{O}}

The hydrochloric acid is a strong acid. It dissociates completely because of weak HCl{\rm{H - Cl}} bond strength. Hence, the mixture of HCl{\rm{HCl}} and NaCl{\rm{NaCl}} cannot act as a buffer.

HCl+NaOHNaCl+H2O{\rm{HCl}} + {\rm{NaOH}} \to {\rm{NaCl}} + {{\rm{H}}_2}{\rm{O}}

The hydrogen iodine is a strong acid because the size of iodide ion very big and the bond length between hydrogen and iodide is very large which causes weak bond strength. Hence, the mixture of HI{\rm{HI}} and NaI{\rm{NaI}} cannot act as a buffer as it dissociates completely.

HI+NaOHNaI+H2O{\rm{HI}} + {\rm{NaOH}} \to {\rm{NaI}} + {{\rm{H}}_2}{\rm{O}}

The hydrogen fluoride is a weak acid because the electronegativity of fluorine is very high which makes the HF{\rm{H - F}} bond difficult to break. It partially dissociates and hence acts a buffer with its conjugate base F{{\rm{F}}^ - } .

HF+NaOHNaF+H2O(Weakacid)(Conjugatebase)\begin{array}{l}\\{\rm{HF }} + {\rm{ NaOH}} \to {\rm{NaF }} + {\rm{ }}{{\rm{H}}_2}{\rm{O}}\\\\\left( \begin{array}{l}\\{\rm{Weak}}\\\\{\rm{acid}}\\\end{array} \right){\rm{ }}\left( \begin{array}{l}\\{\rm{Conjugate }}\\\\{\rm{base}}\\\end{array} \right){\rm{ }}\\\end{array}

Ans:

The mixture of HF{\bf{HF}} and NaF{\bf{NaF}} act as a buffer solution.